OpenStudy (anonymous):

Medal Three barges carry 24 tonnes of gravel each, and a fourth carries 34 tonnes. What is the value, to the nearest dollar, of the whole shipment, at $14.65 per tonne?

OpenStudy (wolf1728):

Total weight = 3*24 plus 34 tons get the total then multiply by 14.65

OpenStudy (wolf1728):

3*24 = 72 plus 34 tons = 106 tons 106 * 14.65 = ?

OpenStudy (anonymous):

1552.9

OpenStudy (wolf1728):

You need value to the nearest dollar so the total is ?

OpenStudy (anonymous):

$1553

OpenStudy (wolf1728):

That's it !!

OpenStudy (anonymous):

so that is the cost of each tonne?

OpenStudy (wolf1728):

No the cost of EACH ton is $14.65

OpenStudy (anonymous):

ok so the answer is $1553

OpenStudy (wolf1728):

That is it

OpenStudy (anonymous):

thank you so much

OpenStudy (wolf1728):

u r welcome